OpenStudy (mhchen):

@triciaal @Directrix @zepdrix @mathmate A sack of 10 balls has 1 red ball. I choose 3 balls. What is the probability that my 3 balls are not red?

OpenStudy (mhchen):

i know this is a hyper-geometric distribution. where n = sample size k=number of successes N= population size x = ....actually idk but I'm guessing trials? :o \[P(X=x) = \frac{\left(\begin{matrix}k \\ x\end{matrix}\right)\left(\begin{matrix}N-k \\ n-x\end{matrix}\right)}{\left(\begin{matrix}N \\ n\end{matrix}\right)}\]

OpenStudy (mhchen):

oh, apparently x = 1 since we need 1 red ball.

OpenStudy (welshfella):

My stats knowledge is pretty rusty but isnt the probaboiity 9/10 * 8/9 * 7/8 ?

OpenStudy (triciaal):

alternate view

OpenStudy (mhchen):

@welshfella The answer is 0.271 btw. I don't think we're using the method u used.

OpenStudy (welshfella):

Oh Ok - then my method is not correct

OpenStudy (mhchen):

Fyi, this is what I thought: \[P(X=x) = \frac{\left(\begin{matrix}1 \\ 1\end{matrix}\right)\left(\begin{matrix}9 \\ 2\end{matrix}\right)}{\left(\begin{matrix}10 \\ 3\end{matrix}\right)}\] Since that finds the probability of getting 1 red ball. (Which turns out to be 0.3)

OpenStudy (mhchen):

and the complement would be...like welsh said..0.7 Triciaal if you get the same answer, i'll just ask my teacher on whether there was a mistake.

OpenStudy (triciaal):